OpenStudy (anonymous):

can anyone pzl help me with math in 8th grade anyone older

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@watchmath

OpenStudy (sleepyjess):

Do you know how to find slope?

OpenStudy (anonymous):

a lil bit

OpenStudy (sleepyjess):

Ok, for the first one, we need to put it in slope-intercept form which is y=mx+b

OpenStudy (anonymous):

wait i think 1 is A positive right

OpenStudy (sleepyjess):

How are you getting that?

OpenStudy (anonymous):

by going to the slope

OpenStudy (sleepyjess):

It isn't in slope intercept form though.

OpenStudy (anonymous):

and number 2 is negative correct

OpenStudy (anonymous):

look it up on google put things in slope form

OpenStudy (sleepyjess):

Number 2 is in slope intercept form, the coefficient before x is the slope.

OpenStudy (anonymous):

o lol so number 2 is A positive cause it shows no negative right

OpenStudy (sleepyjess):

Yes

OpenStudy (sleepyjess):

Now, let's work through number 1 together.

OpenStudy (anonymous):

ik number 3 is negative

OpenStudy (sleepyjess):

2x+3y=-9 Subtract 2x from each side

OpenStudy (sleepyjess):

Yep, number 3 is negative

OpenStudy (anonymous):

i said number is positive

OpenStudy (anonymous):

number 1

OpenStudy (sleepyjess):

But number 1 is not positive

OpenStudy (texaschic101):

for 1 and 2 you will need them in y = mx + b form, where m is the slope. for 3 and 4, you can use slope formula : slope = (y - y1) / (x - x1)

OpenStudy (anonymous):

yea thats wat i meant really lol

OpenStudy (sleepyjess):

You need to understand why number 1 is not positive.

OpenStudy (anonymous):

ok

OpenStudy (sleepyjess):

Now, subtract 2x from each side

OpenStudy (anonymous):